if length 'L',force 'F' and time 'T' are taken as fundamental quantities.What would the dimensional equation of mass and density?
1
Expert's answer
2011-10-14T12:06:17-0400
According to the Second Newton's Law:
we know that F = ma = m*L/T^2.
So, m
= (F*T^2)/L.
Density = m/V = ((F*T^2)/L)/L^3 = (F*T^2)/L^4